# Finetuning Large Language Models for Vulnerability Detection

### v4

@@ 5.3. Imbalanced classification improvement
Focal loss (Lin et al., 2017) applies a modulating factor $(1-p_{t})^{\gamma}$ to the standard cross entropy loss. This factor down-weights well-classified examples and focuses training on hard misclassified cases. The standard cross-entropy loss corresponds to $\gamma=0$.
@@
Table 8. Resulting quality of WizardCoder for different $\gamma$ values in focal loss on the imbalanced $X_{1}$ with $P_{3}$ dataset.
@@
Sample weighting assigns higher importance weights to the $P_{1}$ and $P_{2}$ examples compared to $P_{3}$ during training. Weights from $1.0$ to $30.0$ were tried. Larger weights place more emphasis on the minority informative data. The similar technique was reported to be effective with the focal loss in the original article (Lin et al., 2017).
@@
Table 9. Resulting quality of WizardCoder for different values of $P_{1}+P_{2}$ weights on the imbalanced $X_{1}$ with $P_{3}$ dataset.
@@
Both the focal loss and additional weights do similar work. In the original work (Lin et al., 2017), it was stated that adding weights to the focal loss leads to some improvements.
@@
Table 10. The results of WizardCoder using the focal loss and the sample weighting on the imbalanced $X_{1}$ with $P_{3}$ dataset.
@@ 7. Conclusion
  • A new more precise vulnerability benchmark dataset is collected. It is smaller than most of the available datasets, as well as possesses quality labels that are free from the errors stemming from irrelevant code and cleanup changes (Croft et al., 2023).
@@
While we show an improvement in quality of the vulnerability detection task, the improvement is not so big as we could expect knowing the used LLM’s scale and capacity. This gives us the answer to the question posed at the beginning of the paper: whether the encountered performance limit is due to the limited capacity of CodeBERT-like models. We see that the main bottlenecks of the task that limit the performance lie somewhere else, possibly in the field of dataset quality and usage of the project-level context information.

### v5

@@ 6.3 Imbalanced classification improvement
Focal loss Lin et al. (2017) applies a modulating factor $(1-p_{t})^{\gamma}$ to the standard cross entropy loss. This factor down-weights well-classified examples and focuses training on hard misclassified cases. The standard cross-entropy loss corresponds to $\gamma=0$.
@@
Table 8: Resulting quality of WizardCoder for different $\gamma$ values in focal loss on the imbalanced $X_{1}$ with $P_{3}$ dataset.
@@
Sample weighting assigns higher importance weights to the $P_{1}$ and $P_{2}$ examples compared to $P_{3}$ during training. Weights from $1.0$ to $30.0$ were tried. Larger weights place more emphasis on the minority informative data. The similar technique was reported to be effective with the focal loss in the original article Lin et al. (2017).
@@
Table 9: Resulting quality of WizardCoder for different values of $P_{1}+P_{2}$ weights on the imbalanced $X_{1}$ with $P_{3}$ dataset.
@@
Both the focal loss and additional weights do similar work. In the original work Lin et al. (2017), it was stated that adding weights to the focal loss leads to some improvements.
@@
Table 10: The results of WizardCoder using the focal loss and the sample weighting on the imbalanced $X_{1}$ with $P_{3}$ dataset.
@@ 8 Conclusion
  • A new more precise vulnerability benchmark dataset is collected. It is smaller than most of the available datasets, as well as possesses quality labels that are free from the errors stemming from irrelevant code and cleanup changes Croft et al. (2023).
@@
We show an improvement in quality of the vulnerability detection task by finetuning WizardCoder LLM model for the vulnerability detection task. Considering the quality needed for business tasks, there is still a gap between the achieved and required quality. This leaves us with a big room for improvement, particularly, in choosing the optimal representation of vulnerabilities and usage of a broader project-level context information. These are very exciting prospects for future research.
Appendix A Distribution of CWEs
OtherCWE-862CWE-200CWE-863CWE-352CWE-89CWE-611CWE-79CWE-22CWE-776CWE-264CWE-20CWE-835CWE-295CWE-287CWE-203CWE-755CWE- [^4] CWE-119100100100100200200200200300300300300400400400400500500500500600600600600700700700700800800800800900900900900902902902902504504504504429429429429391391391391324324324324273273273273261261261261201201201201194194194194175175175175153153153153142142142142125125125125919191919090909085858585818181818181818174747474Frequency
Figure 1: Distribution of the top 19 CWE categories and others in the test set with $P_{3}$.
OtherCWE-79CWE-611CWE-20CWE-22CWE-264CWE-502CWE-200CWE-94CWE-863CWE-287CWE-74CWE-78CWE-284CWE-276CWE-755CWE-89CWE-269CWE-352101010102020202030303030404040405050505060606060707070707171717136363636232323232222222213131313101010109999888888888888666666666666666644444444444444444444Frequency
Figure 2: Distribution of the top 19 CWE categories and others in the test set without $P_{3}$.